# Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations

### v1

@@ Abstract
The burgeoning prominence of role-playing agents (RPAs) with large language models (LLMs) is driven by their capacity to model complex social interactions and human-like behaviors in controlled, reproducible environments. However, collecting utterances from a specific character or individual is costly and presents substantial challenges for continuously updating character attributes and model parameters, especially for daily streamers or VTubers whose persona information frequently changes. Although retrieval-augmented generation (RAG) can alleviate this problem, there are three underlying limitations. (i) Previous RAG methods truncate each character’s persona paragraph to a fixed length, regardless of hierarchical contexts and varying lengths across characters, which results in responses with lower persona consistency. (ii) If a persona does not contain knowledge relevant to a given query, RAG-based RPAs are prone to hallucination, making it challenging to generate accurate responses. (iii) Existing role-playing datasets are composed of dialogues, and there is no dataset designed for efficiently constructing and evaluating RPAs based on RAG. To address the aforementioned challenges, we propose Amadeus, which is composed of Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE). ACTS finds an optimal chunk length and hierarchical contexts for each character. AE identifies a character’s general attributes from the chunks retrieved by GS and uses these attributes as a final context to maintain robust persona consistency even when answering out-of-knowledge questions. To facilitate the development and evaluation of RAG-based RPAs, we construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs. We find that our framework effectively models not only the knowledge possessed by characters, but also various attributes such as personality.
Introduction
Large language models with long-context capabilities are engineered to manage lengthy input sequences, allowing them to interpret and utilize extended contextual information. (OpenAI 2025; Qwen et al. 2025; Gemini Team 2025, 2024). Although LLMs exhibit enhanced abilities in understanding extended contexts, they still face significant challenges when handling tasks involving genuinely long contexts (Li et al. 2024a). Furthermore, utilizing all relevant information from long-context models to answer each query can be computationally expensive (Li et al. 2024b).
@@ Introduction
Retrieval-augmented generation (RAG) cost-efficiently mitigates factual inaccuracies and hallucinations in responding to knowledge-intensive queries by integrating external retrieval mechanisms that provide accurate and up-to-date supporting information (Gao et al. 2023; Huang et al. 2025). However, despite these advantages of RAG, there has been little research on RAG-based role-playing agents (RPAs). Moreover, existing role-playing datasets are composed exclusively of dialogues involving characters that are difficult to collect, and there is no dataset designed for building and evaluating RAG-based RPAs.
In this paper, we examine the challenges inherent in RAG-based role-playing and propose approaches to mitigate these challenges. In real-world applications, users and RPAs frequently engage in conversations on topics that extend beyond the knowledge defined in the character’s persona. However, we observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge (Figure 1). To address this issue and maintain persona consistency, we first introduce an Adaptive Context-aware Text Splitter (ACTS), which segments each character’s persona using optimal chunk length and overlap, thereby preserving context across chunks while simultaneously incorporating hierarchical context information within each chunk. Then, the Guided Selection (GS) retrieves appropriate chunks in order to infer information relevant to the question from the character’s persona such as previous actions and behaviors. Finally, the Attribute Extractor (AE) identifies the general attributes of the character from the retrieved chunks, encouraging the RPA to respond in a manner consistent with that character. Note that if GS fails to find a chunk corresponding to the question, AE extracts the character’s attributes using the chunks with the highest similarity scores, which enables consistent responses.
We conduct extensive experiments to examine the factors influencing the performance of RAG-based role-playing, utilizing various settings with 15 fictional character RPAs, the manually constructed CharacterRAG dataset, and multiple psychological questionnaire datasets. Results demonstrate that our framework reveals new possibilities for RAG-based RPAs. We summarize our contributions as follows:
  • We propose AMADEUS, a RAG-based RPA framework that not only elicits information related to a character, but also maintains persona consistency even when responding to queries beyond its explicit knowledge.
@@
  • We manually construct CharacterRAG, a role-playing dataset for implementing and evaluating RAG-based RPAs comprising persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
  • We systematically investigate and uncover key considerations for building RAG-based RPAs through extensive experiments performed in a range of settings.

### v3

@@ Abstract
Building role-playing agents (RPAs) that faithfully emulate specific characters remains challenging because collecting character-specific utterances and continually updating model parameters are resource-intensive, making retrieval-augmented generation (RAG) a practical necessity. However, despite the importance of RAG, there has been little research on RAG-based RPAs. For example, we empirically find that when a persona lacks knowledge relevant to a given query, RAG-based RPAs are prone to hallucination, making it challenging to generate accurate responses. In this paper, we propose Amadeus, a training-free framework that can significantly enhance persona consistency even when responding to questions that lie beyond a character’s knowledge. In addition, to underpin the development and rigorous evaluation of RAG-based RPAs, we manually construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs. We find that our proposed method effectively models not only the knowledge possessed by characters, but also various attributes such as personality. The code and dataset will be made publicly available at our GitHub.
Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations
Jeiyoon Park1, Yongshin Han1, Minseop Kim1, Kisu Yang2, 1SOOP, 2Korea University {naruto, winterfeb, usio}@sooplive.com, willow4@korea.ac.kr https://huggingface.co/datasets/naruto-soop/CharacterRAG
Figure 1: Comparison of Responses between Naive RAG and Our Method. We observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge.
1 Introduction
@@
Large language models with long-context capabilities are engineered to manage lengthy input sequences, allowing them to interpret and utilize extended contextual information (OpenAI, 2025; Qwen et al., 2025; Gemini Team, 2025, 2024). Although LLMs exhibit enhanced abilities in understanding extended contexts, they still face significant challenges when handling tasks involving genuinely long contexts (Li et al., 2024a). Furthermore, utilizing all relevant information from long-context models to answer each query can be computationally expensive (Li et al., 2024b). Retrieval-augmented generation (RAG) cost-efficiently mitigates factual inaccuracies and hallucinations in responding to knowledge-intensive queries by integrating external retrieval mechanisms that provide accurate and up-to-date supporting information (Gao et al., 2023; Huang et al., 2025). However, despite these advantages of RAG, there has been little research on RAG-based role-playing agents (RPAs). Moreover, existing role-playing datasets are composed exclusively of dialogues involving characters that are difficult to collect, and there is no dataset designed for building and evaluating RAG-based RPAs. We examine the challenges inherent in RAG-based role-playing and propose approaches. In real-world applications, users and RPAs frequently engage in conversations on topics that extend beyond the knowledge defined in the character’s persona. However, we observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge (Figure 1 and Figure 6).
To address these challenges, we introduce Amadeus, a training-free framework that can markedly improves persona consistency, even when addressing questions beyond a character’s knowledge. Amadeus consists of three substages: Adaptive Context-aware Text Splitter (ACTS), which segments a persona for role-playing, Guided Selection (GS), which retrieves appropriate chunks to infer information relevant to the question from the persona, such as prior actions and behaviors, and Attribute Extractor (AE), which identifies general attributes of the character from the retrieved chunks, thereby encouraging the RPA to respond in a manner consistent with that character.
To underpin the development and rigorous evaluation of RAG-based RPAs, we manually construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs. We conduct extensive experiments to examine the factors that influence the performance of RAG-based RPAs through interview-based assessments (Wang et al., 2024b; Park et al., 2025; Jiang et al., 2023). To investigate whether RPAs can respond appropriately to questions that fall outside their endowed knowledge, we employ multiple psychological questionnaires such as MBTI [^1] https://www.16personalities.com/ and BFI (Barrick and Mount, 1991), following prior studies Sang et al. (2022); Wang et al. (2024b). Results show that our framework opens up new possibilities for RAG-based role-playing.
2 CharacterRAG
2.1 Dataset Construction
We construct the CharacterRAG dataset, which consists of 15 fictional characters, to leverage and evaluate a RAG-based role-playing framework. CharacterRAG is a high-quality, role-playing dataset in which all external information about works featuring characters that could affect persona consistency has been manually removed, and each persona document has been directly reconstructed from the perspective of each character by human annotators [^2] CharacterRAG dataset is sourced from Namuwiki and is based on Korean data: https://namu.wiki/. For instance, any information speculated from the perspective of editors rather than the characters themselves, as well as information such as character popularity polls that may disrupt role-playing, is excluded. CharacterRAG consists of 15 distinct fictional characters, 976K written characters, and 450 question–answer pairs.
@@ 2.1 Dataset Construction
Figure 2: An overview of CharacterRAG Dataset. CharacterRAG consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
2.2 Attributes
Six commonly used attributes in role‑playing define each character’s persona and the corresponding question–answer pairs (Chen et al., 2025): (1) Activity. A documented history comprising prior activities, behaviors, and interactions, encompassing elements such as backstory and schedules. (2) Belief and Value. The foundational tenets, dispositions, and ideological orientations that inform and guide a character’s viewpoints and decision-making processes (e.g., beliefs and attitudes). (3) Demographic Information. Information that can identify a character, including but not limited to their name, age, educational background, professional history, and geographic location. (4) Psychological Traits. Attributes associated with personality traits, emotional states, preferences, and patterns of cognitive behavior. (5) Skill and Expertise. The extent of understanding, skillfulness, and competence regarding particular fields or technologies. (6) Social Relationships. The characteristics and processes of social interactions, encompassing individuals’ roles, relational ties, and patterns of communication. Each section of the character’s persona contains subsections, preserving the hierarchical information (e.g., "Tanjiro Kamado’s actions in the story"). Furthermore, Each QA pair consists of a question and corresponding answer derived from the character’s knowledge, pertaining to one of the six attributes manually constructed for each character. Statistics and samples of CharacterRAG are shown in Figure 2 and Table 7.
3 Task Formulation: RAG-based Role-Playing Agents
Given user query $u$, RAG-based RPAs can be formulated as: $\mathcal{R}=f(u,\mathcal{D}_{p})$, where $\mathcal{D}_{p}$ is a character’s persona, and $f$ is a RPA. A text splitter $g$ divides the persona into $n$ chunks as follows:
$g(\mathcal{D}_{p})=\{c_{1},c_{2},...,c_{n}\}$ (1)
Each chunk $c_{i}$ contains a character’s knowledge corresponding to attributes in Section 2.2. Rather than using the full chunks $\mathcal{C}=\{c_{i}\mid i=1,\dots,n\}$, $f$ takes as input the top $K$ chunks with the highest semantic scores relative to the $u$:
$\mathcal{C}^{*}=\text{TopK}(\{\text{sim}(u,c_{i})\}_{i=1}^{n}),\quad|\mathcal{C}^{*}|=K$ (2)
The objective of $f$ is to vividly embody a character and generate response $\mathcal{R}$ to $u$ while maintaining persona consistency: $\mathcal{R}^{*}=f(u,\mathcal{C}^{*})$. However, previous RAG methods (Guu et al., 2020; Guo et al., 2024; Yang et al., 2024; Shukla et al., 2025; Wang et al., 2025b) truncate each character’s persona paragraph to a fixed length, regardless of the varying lengths across characters, which results in hallucinations or responses with lower persona consistency. Although existing works (LangChain, 2023; Antematter, 2024; Zhong et al., 2025a; Liu et al., 2025) explore optimal chunking strategies, they struggle to capture the contextual similarities across chunks that are crucial for role-playing.
4 Method
Amadeus consists of three substages: Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), in order to build realistic RAG-based RPAs (Figure 5).
4.1 Adaptive Context-aware Text Splitter
Unlike previous naive semantic chunking methods, Given $\mathcal{D}_{p}$, ACTS aims to preserve intra-level context across chunks and, for each chunk, information about the corresponding subsections of the persona—that is, hierarchical context $\mathcal{H}$. For instance, in Figure 2, chunks within ### 4.1.1 must preserve $\mathcal{H}$: "Tanjiro Kamado’s actions in the story (# 4)", "The death of his family (## 4.1)", and "Tanjiro vs. the Water Hashira, Giyu Tomiok (### 4.1.1)".
To this end, ACTS first finds the maximum length of the paragraphs that constitute the persona: $l_{\text{max}}=\varphi(p_{1},p_{2},...,p_{l})$, where $\varphi$ denotes a length-calculating function. Then, ACTS sets the overlap length of the text splitter to half of $l_{\text{max}}$ (i.e., $l_{\text{o}}=l_{\text{max}}/2$). The reason for setting both chunk length and overlap length sufficiently large is to minimize information loss, as the context between pieces of information contained in each chunk is indispensable in RAG-based role-playing (Figure 8). Finally, ACTS recursively retrieves the context at each chunk’s current position in the hierarchy, then segments $\mathcal{D}_{p}$ using $l_{\text{max}}$ and $l_{\text{o}}$, and concatenates the resulting context $\mathcal{H}_{i}$ to each chunk to preserve information such as character descriptions and situational context at each point in the narrative: $\text{ACTS}(\mathcal{D}_{p},\mathcal{H},l_{\text{max}},l_{\text{o}})=\{\hat{c}_{1},\hat{c}_{2},...,\hat{c}_{m}\}$, where $\hat{c}_{i}=[c_{i};\mathcal{H}_{i}]$. From a computational standpoint, the extraction of hierarchical context incurs an O(N) runtime cost, making it efficient.
4.2 Guided Selection
GS focuses on selecting appropriate chunks to generate natural and persona-consistent responses. GS is composed of three stages: (1) We iterate over the chunks, which are sorted in descending order of semantic similarity to the user query $u$, and employ an LLM to determine whether it is possible to infer the corresponding character’s attributes from each chunk for the $u$. (2) The chunks selected in the previous step are appended to the slot, and the iteration terminates when the slot is full. (3) If the slot remains empty after the maximum number of search iterations, the $K$ chunks with the highest semantic similarity to the query is returned. Note that GS is effective in identifying chunks containing information that can be inferred from a character’s actions, such as beliefs or personality traits, which are not explicitly stated in the knowledge base and therefore are difficult to retrieve through direct search. For example, even if there is no explicit knowledge corresponding to the query "My living and work spaces are clean and organized", if Megumin’s conscientiousness can be inferred from her behavior depicted in the narrative, RPA can leverage this information to infer the characteristics of Megumin and generate an appropriate response. GS is summarized in Algorithm 1 and Figure 3.
4.3 Attribute Extractor
Inspired by the observations that incorporating character’s attributes can lead to more realistic responses (Park et al., 2025; Chen et al., 2025), AE considers two attributes: Belief and Value, and Psychological Traits. Beliefs and values are fundamental principles and ideological orientations that inform and influence a character’s viewpoints and choices. On the other hand, psychological traits refer to characteristics related to personality, emotional states, personal interests, and cognitive tendencies. AE extracts attributes of a character from the chunks generated as a result of GS, and exploits them as context. Finally, we dynamically construct the context using knowledge retrieved via RAG and attributes extracted by GS and AE, enabling the model to generate vivid responses (Figure 7).
5 Experiments
5.1 Setup